40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14527     Entry time: Tue Apr 9 18:44:00 2019     In reply to: 14524
Author: gautam 
Type: Update 
Category: ALS 
Subject: EX Green PDH discriminant measurement 

I decided to use the more direct method, of disconnecting feedback to the EX laser PZT, and then looking at the cavity flashes. 

Attachment #1 shows the cavity swinging through two resonances (data collected via oscilloscope). Traces are for the demodulated PDH error signal (top) and the direct photodiode signal (bottom). The traces don't look very clean - I wonder if some saturation / slew rate effects are at play, because we are operating the PD in the 30 dB setting, where the bandwidth of the PD is spec-ed as 260 kHz, whereas the dominant frequency component of the light on the PD is 430 kHz.

The asymmetric horns corresponding to the sideband resonances were also puzzling. Doing the modeling, Attachment #2, I think this is due to the fact that the demodulation phase is poorly set. The PDH modulation frequency is only ~5x the cavity linewidth, so both the real and imaginary parts of the cavity reflectivity contribute to the error signal. If this calculation is correct, we can benefit (i.e. get a larger PDH discriminant) by changing the demod phase by 60 degrees. However, for 230 kHz, it is impractical to do this by just increasing cable length between the function generator and mixer.

Anyway, assuming that we are at the phi=30 degree situation (since the measurement shows all 3 horns going through roughly the same voltage swing), the PDH discriminant is ~40 uV/Hz. In lock, I estimate that there is ~60 uW of light incident on the PDH reflection photodiode. Using the PD response of 0.2 A/W, transimpedance of 47.5 kohm, and mixer conversion loss of 6dB, the shot-noise limited sensitivity is 0.5 mHz/rtHz. The photodiode dark noise contribution is a little lower - estimated to be 0.2 mHz/rtHz. The loop does not have enough gain to reach these levels.

Quote:

PDH discriminant (40 uV/Hz, see this elog) 

Attachment 1: cavityFlashes.pdf  198 kB  Uploaded Tue Apr 9 19:56:16 2019  | Show | Hide all | Show all
Attachment 2: modelPDH.pdf  129 kB  Uploaded Tue Apr 9 19:56:27 2019  | Hide | Hide all | Show all
modelPDH.pdf
ELOG V3.1.3-